OpenStudy (anonymous):

How many solutions can be found for the equation 4x + 5 = 10? Zero One Two Infinitely Many

OpenStudy (alexandervonhumboldt2):

solve: 4x+5=10 4x+5-5=10-5 4x=5 x=5/4. as you can see there is only 1 solution.

OpenStudy (chaise):

One solution, that solution is x. You can solve for x by taking away 5 from both sides, then dividing by 4.

OpenStudy (anonymous):

thanks for the help!!!!

OpenStudy (alexandervonhumboldt2):

np